# Aioli: A unified optimization framework for language model data mixing

### v1

@@ C.2 Values of mixing law parameters
We explain how to compare method-specific $A^{t}$’s to an approximation of the true $A^{t\star}$. First, after performing method-specific initialization, such as training reference models, we run each online method (Skill-It, DoReMi’s proxy model DoGE’s proxy model, Skill-it, and Aioli) for $t$ steps. For Skill-It, DoReMi, and DoGE, we use the unrestricted setting configuration of hyperparameters presented in Section D. For Aioli, we analyze the parameters of Aioli +GS from the restricted setting, since we found that this had less noisy fluctuation in the weights than in the unrestricted setting. For $m=2$, we set $t=1000$ for Skill-It and $t=500$ for DoGE, DoReMi, and Aioli since Skill-It is updated less frequently. For $m=3$, we set $t=1000$ for DoGE, DoReMi and Skill-It, and $t=1500$ for Aioli. We then checkpoint the language model and the method’s $A^{t}$. For DoGE and DoReMi, we compute a smoothed $A^{t}=\frac{1}{100}\sum_{i=1}^{100}A^{t-100+i}$ because each $A^{t}$ is computed at the batch level, and can thus be noisy. For Aioli, we also smooth the $A^{t}$ by averaging the previous timestep parameters.
To approximate $A^{t\star}$, we run a training sweep over $\mathcal{P}$ for the next 100 steps after the checkpoint. We use this training sweep to fit $A^{t\star}$ from the dynamic mixing law $L_{\text{val},i}^{t+1}(\bm{p})=L_{\text{val},i}^{t}(\bm{p})-\sum_{j=1}^{m}A_{ij}^{t\star}p_{j}^{t}$.
Before we compare parameters, we must scale $A^{t}$ by some $b^{t}$ where $L_{\text{val},i}^{t+1}(\bm{p})=L_{\text{val},i}^{t}(\bm{p})-b^{t}\sum_{j=1}^{m}A_{ij}^{t}p_{j}^{t}$ for all $i\in[m]$. This is allowed since $b^{t}$ does not influence the optimal $\bm{p}$ and does not need to be in the update rule. We fit a single $b^{t}$ across all groups’ mixing laws and set $\tilde{A}^{t}=b^{t}A^{t}$. We can then compare $A^{t}$ and $A^{t\star}$ using the metric $\text{sim}(\tilde{A}^{t},A^{t\star})=0.5\text{cossim}(\tilde{a}^{t},a^{t\star})+0.5\text{Spearman}(\tilde{a}^{t},a^{t\star})$, which we proposed in Section 4.3.
@@ C.2.1 Properties of $A^{t\star}$
We discuss some properties of $A^{t\star}$, finding that 1) $A^{t\star}$ can vary significantly across time, and 2) $A^{t\star}$ needs to be modeled as a full matrix. In Appendix E.2, we also present ablations on Aioli that test how capturing these two properties in Aioli are important to performance.
For each initial mixture $p^{0}\in\mathcal{P}$, we train for $t=2000$ steps and then sweep over $\mathcal{P}$ for the next $100$ steps. We repeat this setup for $t=4000$ to obtain $A^{2000\star}$ and $A^{4000\star}$. We do this experiment for Arxiv/Stackexchange and Github/C4.
@@
However, for Arxiv/Stackexchange, we found that the column sums of $A^{2000\star}$ and $A^{4000\star}$ never change in terms of the ordering of proportions of the data groups, across all $p^{0}\in\mathcal{P}$. As a result, the optimal $p^{t}$ never changes direction. This suggests that how much $A^{t}$ varies in ordering over time depends on the data groups. As a result, methods like Skill-It, which use a time-invariant $A^{\text{SG}}$ multiplied by validation loss, may not be able to match the true $A^{t\star}$ if the groups’ validation losses change in ranking across time, which we observe in Github/C4.
Modeling $A^{t\star}$ as a full vs diagonal matrix. We find that modeling the off-diagonal entries of $A^{t\star}$ is important. For each sweep, we fit both $A^{t\star}$ as described above and a diagonal matrix $A_{d}^{t\star}$. We compare if the column sums of $A^{t\star}$ and $A_{d}^{t\star}$ differ in the order of elements.
@@
Using the diagonal matrix for Github/C4 would result in prioritizing training on Github, even though the full matrix suggests that C4 should be prioritized. Therefore, it is important to model $A^{t\star}$ as a full matrix. Methods like DoReMi, which use a diagonal $A^{t}$, can perform suboptimally.
@@ C.3 Solving strategy
Figure 5: The linear dynamic parameterization results from Figure 2, with $p^{t}=[0,1]$ and $[1,0]$ also plotted. We see that the linear dynamics are misspecified at $p_{i}^{t}=0$ for both $i$.
@@
Appendix D Experimental Details
Training details.
@@ Appendix D Experimental Details
To obtain the SlimPajama test sets, we shuffle and split the validation set from SlimPajama-6B [53, 73] in half. Next, we discuss the training setups for the restricted and unrestricted settings. For the $m=2,3$ settings, we train a 160M model using Pythia-160M’s configuration for $S=5000$ steps and results are averaged over 5 random seeds. For $m=7$, we train a 160M model using Pythia-160M’s configuration for $S=40000$ steps results are averaged over 3 random seeds. All settings use FlashAttention [17], batch size of 8, context size of 2048, and cosine learning rate decay from a starting learning rate of 5e-5 to 1e-5 with 500 steps of learning rate warmup.
@@
Mixing algorithm details.
See Table 12 for a glossary of hyperparameters. $m=2$: restricted Table 13, unrestricted Table 14; $m=3$: restricted Table 15, unrestricted Table 16; $m=7$: restricted Table 17, unrestricted Table 18.
In Table 19, we provide the mixture proportions for each method (averaged across training steps) for each dataset on one random seed. In Figure 6, we provide all of Aioli’s proportion trajectories throughout training in both the unrestricted and restricted settings on one random seed for the $m=2$ settings.
[…]
@@
In the unrestricted setting, we found it sometimes helpful to use an exponential moving average with proportion $\gamma$ over $A^{t}$ for Aioli. Formally, the standard $p^{t}$ update rule in Algorithm 1 can be unrolled as $p_{j}^{t+1}\propto p_{j}^{0}\exp(\eta\sum_{\tau=1}^{t}\sum_{i=1}^{m}\bar{A}_{ij}^{\tau})$, which places equal weight on every $\bar{A}_{ij}^{\tau}$. To incorporate the EMA, we define $A^{1}_{\text{ema}}=\bar{A}^{1}$ and $A^{t}_{\text{ema}}=(1-\gamma)\bar{A}^{t}+\gamma A^{t-1}_{\text{ema}}$. We then use the update rule $p_{j}^{t+1}\propto p_{j}^{0}\exp(\eta A_{\text{ema}}^{t})$. This allows Aioli to decay the contributions of $A^{t}$, such that the value of $p^{t}$ is less dependent on earlier proportions in the training.
Table 12: Hyperparameters in each of the online mixing methods. Hyperparameters for grid search and DML are the number of runs and number of steps per run, which are described above.
Method Hyperparameter Type
Aioli $\cdot$ number of rounds $T$ int
$\cdot$ sweeps $k$ int
$\cdot$ proportion of round $\delta$ dedicated to LearnParams float
$\cdot$ EGD learning rate $\eta$ float
$\cdot$ $\varepsilon$ one-hot smoothing factor float
$\cdot$ EMA parameter $\gamma$ float or None
Skill-it $\cdot$ number of rounds $T$ int
$\cdot$ EGD learning rate $\eta$ float
$\cdot$ multiplicative weights window int
DoReMi $\cdot$ EGD learning rate $\eta$ float
DoGE $\cdot$ EGD learning rate $\eta$ float
$\cdot$ proportion of training batch that consists of validation set. DoGE computes gradients on the validation set float
@@
Data groups Method Hyperparameter Value
arXiv/SE Aioli $\cdot$ number of rounds $T$ 20
$\cdot$ sweeps $k$ 4
$\cdot$ proportion of round $\delta$ 0.128
$\cdot$ $\varepsilon$ one-hot smoothing factor 0.75
@@
Skill-it $\cdot$ number of rounds $T$ 10
$\cdot$ EGD learning rate $\eta$ 0.2
$\cdot$ Multiplicative weights window. 3
DoReMi $\cdot$ EGD learning rate $\eta$ 0.01
DoGE $\cdot$ EGD learning rate $\eta$ 0.01
$\cdot$ proportion of batch for validation set 0.25
GitHub/C4 Aioli $\cdot$ number of rounds $T$ 20
$\cdot$ sweeps $k$ 4
$\cdot$ proportion of round $\delta$ 0.128
$\cdot$ $\varepsilon$ one-hot smoothing factor 0.75
@@
Skill-it $\cdot$ number of rounds $T$ 10
$\cdot$ EGD learning rate $\eta$ 0.2
$\cdot$ Multiplicative weights window. 3
DoReMi $\cdot$ EGD learning rate $\eta$ 0.01
DoGE $\cdot$ EGD learning rate $\eta$ 0.1
$\cdot$ proportion of batch for validation set 0.25
Books/SE Aioli $\cdot$ number of rounds $T$ 20
$\cdot$ sweeps $k$ 4
$\cdot$ proportion of round $\delta$ 0.128
$\cdot$ $\varepsilon$ one-hot smoothing factor 0.75
@@
Skill-it $\cdot$ number of rounds $T$ 10
@@
$\cdot$ Multiplicative weights window. 3
DoReMi $\cdot$ EGD learning rate $\eta$ 0.01
DoGE $\cdot$ EGD learning rate $\eta$ 0.01
$\cdot$ proportion of batch for validation set 0.25
Data groups Method Hyperparameter Value
arXiv/SE Aioli $\cdot$ number of rounds $T$ 20
$\cdot$ sweeps $k$ 4
$\cdot$ proportion of round $\delta$ 0.128
$\cdot$ $\varepsilon$ one-hot smoothing factor 0.75
@@
Skill-it $\cdot$ number of rounds $T$ 10
$\cdot$ EGD learning rate $\eta$ 0.2
$\cdot$ Multiplicative weights window. 3
DoReMi $\cdot$ EGD learning rate $\eta$ 0.01
DoGE $\cdot$ EGD learning rate $\eta$ 0.01
$\cdot$ proportion of batch for validation set 0.25
GitHub/C4 Aioli $\cdot$ number of rounds $T$ 20
$\cdot$ sweeps $k$ 4
$\cdot$ proportion of round $\delta$ 0.128
$\cdot$ $\varepsilon$ one-hot smoothing factor 0.75
$\cdot$ EGD learning rate $\eta$ 0.3
$\cdot$ EMA parameter $\gamma$ 0.5
Skill-it $\cdot$ number of rounds $T$ 5
$\cdot$ EGD learning rate $\eta$ 0.1
$\cdot$ Multiplicative weights window. 3
DoReMi $\cdot$ EGD learning rate $\eta$ 0.01
DoGE $\cdot$ EGD learning rate $\eta$ 0.1
$\cdot$ proportion of batch for validation set 0.25
Books/SE Aioli $\cdot$ number of rounds $T$ 20
$\cdot$ sweeps $k$ 4
$\cdot$ proportion of round $\delta$ 0.128
$\cdot$ $\varepsilon$ one-hot smoothing factor 0.75
$\cdot$ EGD learning rate $\eta$ 0.1
$\cdot$ EMA parameter $\gamma$ None
Skill-it $\cdot$ number of rounds $T$ 10
$\cdot$ EGD learning rate $\eta$ 0.8
$\cdot$ Multiplicative weights window. 3
DoReMi $\cdot$ EGD learning rate $\eta$ 0.01
DoGE $\cdot$ EGD learning rate $\eta$ 0.01
$\cdot$ proportion of batch for validation set 0.25
Data groups Method Hyperparameter Value
arXiv/Books/SE Aioli $\cdot$ number of rounds $T$ 20
$\cdot$ sweeps $k$ 4
$\cdot$ proportion of round $\delta$ 0.288
$\cdot$ $\varepsilon$ one-hot smoothing factor 0.75
$\cdot$ EGD learning rate $\eta$ 0.2
$\cdot$ EMA parameter $\gamma$ None
Skill-it $\cdot$ number of rounds $T$ 10
$\cdot$ EGD learning rate $\eta$ 0.2
$\cdot$ Multiplicative weights window. 3
DoReMi $\cdot$ EGD learning rate $\eta$ 0.01
DoGE $\cdot$ EGD learning rate $\eta$ 0.01
$\cdot$ proportion of batch for validation set 0.5
CommonCrawl/GitHub/Wiki Aioli $\cdot$ number of rounds $T$ 20
$\cdot$ sweeps $k$ 4
$\cdot$ proportion of round $\delta$ 0.288
$\cdot$ $\varepsilon$ one-hot smoothing factor 0.75
$\cdot$ EGD learning rate $\eta$ 0.2
$\cdot$ EMA parameter $\gamma$ None
Skill-it $\cdot$ number of rounds $T$ 10
$\cdot$ EGD learning rate $\eta$ 0.2
$\cdot$ Multiplicative weights window. 3
DoReMi $\cdot$ EGD learning rate $\eta$ 0.01
DoGE $\cdot$ EGD learning rate $\eta$ 0.01
$\cdot$ proportion of batch for validation set 0.5
Data groups Method Hyperparameter Value
arXiv/Books/SE Aioli $\cdot$ number of rounds $T$ 20
$\cdot$ sweeps $k$ 4
$\cdot$ proportion of round $\delta$ 0.288
$\cdot$ $\varepsilon$ one-hot smoothing factor 0.75
$\cdot$ EGD learning rate $\eta$ 0.1
$\cdot$ EMA parameter $\gamma$ 0.5
Skill-it $\cdot$ number of rounds $T$ 10
$\cdot$ EGD learning rate $\eta$ 0.2
$\cdot$ Multiplicative weights window. 3
DoReMi $\cdot$ EGD learning rate $\eta$ 0.01
DoGE $\cdot$ EGD learning rate $\eta$ 0.01
$\cdot$ proportion of batch for validation set 0.5
CommonCrawl/GitHub/Wiki Aioli $\cdot$ number of rounds $T$ 20
$\cdot$ sweeps $k$ 4
$\cdot$ proportion of round $\delta$ 0.288
$\cdot$ $\varepsilon$ one-hot smoothing factor 0.75
$\cdot$ EGD learning rate $\eta$ 0.3
$\cdot$ EMA parameter $\gamma$ 0.5
Skill-it $\cdot$ number of rounds $T$ 10
$\cdot$ EGD learning rate $\eta$ 0.2
$\cdot$ Multiplicative weights window. 3
DoReMi $\cdot$ EGD learning rate $\eta$ 0.01
DoGE $\cdot$ EGD learning rate $\eta$ 0.01
$\cdot$ proportion of batch for validation set 0.5
@@
Data groups Method Hyperparameter Value
SlimPajama, full Aioli $\cdot$ number of rounds $T$ 20
$\cdot$ sweeps $k$ 2
$\cdot$ proportion of round $\delta$ 0.07
$\cdot$ $\varepsilon$ one-hot smoothing factor 0.75

### v2

@@ C.2 Values of mixing law parameters
We explain how to compare method-specific $A^{t}$’s to an approximation of the true $A^{t\star}$. First, after performing method-specific initialization, such as training reference models, we run each online method (Skill-It, DoReMi’s proxy model DoGE’s proxy model, Skill-it, and Aioli) for $t$ steps. For Skill-It, DoReMi, and DoGE, we use the unrestricted setting configuration of hyperparameters presented in Section E. For Aioli, we analyze the parameters of Aioli +GS from the restricted setting, since we found that this had less noisy fluctuation in the weights than in the unrestricted setting. For $m=2$, we set $t=1000$ for Skill-It and $t=500$ for DoGE, DoReMi, and Aioli since Skill-It is updated less frequently. For $m=3$, we set $t=1000$ for DoGE, DoReMi and Skill-It, and $t=1500$ for Aioli. We then checkpoint the language model and the method’s $A^{t}$. For DoGE and DoReMi, we compute a smoothed $A^{t}=\frac{1}{100}\sum_{i=1}^{100}A^{t-100+i}$ because each $A^{t}$ is computed at the batch level, and can thus be noisy. For Aioli, we also smooth the $A^{t}$ by averaging the previous timestep parameters.
To approximate $A^{t\star}$, we then run a training sweep of $p^{t}$ over $\mathcal{P}$ for 100 steps on the checkpoint. We use this training sweep to fit $A^{t\star}$ from the dynamic mixing law $L_{\text{val},i}^{t+1}(\bm{p})=L_{\text{val},i}^{t}(\bm{p})-\sum_{j=1}^{m}A_{ij}^{t\star}p_{j}^{t}$.
Before we compare parameters, we scale $A^{t}$ by some $b^{t}$ where $L_{\text{val},i}^{t+1}(\bm{p})=L_{\text{val},i}^{t}(\bm{p})-b^{t}\sum_{j=1}^{m}A_{ij}^{t}p_{j}^{t}$ for all $i\in[m]$. This is allowed since $b^{t}$ does not influence the optimal $\bm{p}$ and does not need to be in the update rule. We fit a single $b^{t}$ across each group’s mixing law and set $\tilde{A}^{t}=b^{t}A^{t}$. We can then compare $A^{t}$ and $A^{t\star}$ using the metric $\text{sim}(\tilde{A}^{t},A^{t\star})=0.5\text{cossim}(\tilde{a}^{t},a^{t\star})+0.5\text{Spearman}(\tilde{a}^{t},a^{t\star})$, which we proposed in Section 4.3.
@@ C.2.1 Properties of $A^{t\star}$
We discuss some properties of $A^{t\star}$, finding that 1) $A^{t\star}$ can vary significantly across time, and 2) $A^{t\star}$ needs to be modeled as a full matrix. To do this, for each initial mixture $p^{0}\in\mathcal{P}$, we train for $t=2000$ steps and then sweep over $\mathcal{P}$ for the next $100$ steps. We repeat this setup for $t=4000$ to obtain $A^{2000\star}$ and $A^{4000\star}$. We do this experiment for Arxiv/Stackexchange and Github/C4.
@@
However, for Arxiv/Stackexchange, the column sums of $A^{2000\star}$ and $A^{4000\star}$ never change in terms of the ordering of proportions of the data groups, across all $p^{0}\in\mathcal{P}$. As a result, the optimal $p^{t}$ never changes direction. This suggests that how much $A^{t}$ varies in ordering over time depends on the data groups. As a result, methods like Skill-It, which use a time-invariant $A^{\text{SG}}$ multiplied by validation loss, may not be able to match the true $A^{t\star}$ if the groups’ validation losses do not change in ranking across time, which we observe in Github/C4.
Modeling $A^{t\star}$ as a full vs diagonal matrix. We find that modeling the off-diagonal entries of $A^{t,\star}$ is important. For each sweep, we fit both $A^{t\star}$ as described above and a diagonal matrix $A_{d}^{t\star}$. We compare if the column sums of $A^{t\star}$ and $A_{d}^{t\star}$ differ in the order of elements.
@@
Using the diagonal matrix for Github/C4 would result in prioritizing training on Github, even though the full matrix suggests that C4 should be prioritized. Therefore, it is important to model $A^{t\star}$ as a full matrix. As a result, methods like DoReMi, which use a diagonal $A^{t}$, can perform suboptimally.
@@ C.3 Solving strategy
Figure 6: The linear dynamic parameterization results from Figure 2 (right), with $p^{t}=[0,1]$ and $[1,0]$ also plotted. We see that the linear dynamics are misspecified at $p_{i}^{t}=0$ for both $i$.
@@
Appendix D Additional Algorithmic Details
In Aioli, LearnParams is used in each round to learn $A^{t}$. Then, $A^{t}$ is used to compute $p^{t}$, which is used for training during the round. We provide a derivation of LearnParams by first presenting a naive, high-cost method for estimating $A^{t}$ (Appendix D.1). This involves checkpointing the model at each round, running a training sweep over the round and observing the changes in validation losses, and fitting $A^{t}$ to these changes. Then, we layer on two modifications that compute slightly different loss changes, helping lower the cost of estimation. First, we shorten the training sweep to be only over a fraction of the round, $\delta$, and use these shortened changes in validation losses to fit $A^{t}$ (Appendix D.2). Second, we simulate a simultaneous training sweep by partitioning the $\delta$ fraction of the round into many small parts, interleaving the different sweep mixtures at a fine granularity and averaging the loss changes for each sweep mixture (Appendix D.3). This idea, with similarity to concepts like time-division multiplexing in signal processing [12], enables Aioli to require no extra training while trading off accuracy of the estimate. We provide a sketch of our derivation in Figure 7.
Figure 7: Derivation of Aioli. Top: a naive high-cost approach where training sweeps are conducted to fit $A^{t}$ at each round (Appendix D.1). Middle: a modification that shortens the training sweeps used to learn $A^{t}$ (Appendix D.2). Bottom: a final modification that interleaves the sweep mixtures at a high frequency (large $k$) in one single run, enabling Aioli’s LearnParams to require no additional training (Appendix D.3).
D.1 Naive training sweep approach
This approach is depicted in Figure 7 (top). By conducting a training sweep over round $t$, we can use a linear system of equations to estimate $A^{t}$ from the linear dynamic mixing law $L_{\text{val},i}^{t+1}(\bm{p})=L_{\text{val},i}^{t}(\bm{p})-\sum_{j=1}^{m}A_{ij}^{t}p_{j}^{t}$. Let $p^{t,1},p^{t,2},\dots p^{t,m}\in\triangle^{m}$ comprise a training sweep over the duration of round $t$. First, we checkpoint the model $f^{t}$, and for simplicity denote $f^{t}$’s validation loss on group $i$ as $L_{\text{val},i}^{t}$. For each $p^{t,j}$, we train $f^{t}$ for the entire round using $p^{t,j}$. We then record how much the validation loss on each group changes, $L_{\text{val},i}^{t}-L_{\text{val},i}^{t+1}(p^{t,j})$ for all $i\in[m]$. By the end of this procedure on each $p^{t,j}$, we have the following system of equations for each $i\in[m]$:
$\sum_{j=1}^{m}A_{ij}^{t}p_{j}^{t,1}$ $=L_{\text{val},i}^{t}-L_{\text{val},i}^{t+1}(p^{t,1})$ (13)
$\sum_{j=1}^{m}A_{ij}^{t}p_{j}^{t,2}$ $=L_{\text{val},i}^{t}-L_{\text{val},i}^{t+1}(p^{t,2})$
$\vdots$
$\sum_{j=1}^{m}A_{ij}^{t}p_{j}^{t,m}$ $=L_{\text{val},i}^{t}-L_{\text{val},i}^{t+1}(p^{t,m})$
This is a system of linear equations with $m$ unknowns: $A_{i1},\dots,A_{im}$. We can write it in matrix form as:
$\begin{bmatrix}p_{1}^{t,1}&p_{2}^{t,1}&\dots&p_{m}^{t,1}\\ p_{1}^{t,2}&p_{2}^{t,2}&\dots&p_{m}^{t,2}\\ \vdots&&&\\ p_{1}^{t,m}&p_{2}^{t,m}&\dots&p_{m}^{t,m}\end{bmatrix}\begin{bmatrix}A_{i1}^{t}\\ A_{i2}^{t}\\ \vdots\\ A_{im}^{t}\end{bmatrix}=\begin{bmatrix}L_{\text{val},i}^{t}-L_{\text{val},i}^{t+1}(p^{t,1})\\ L_{\text{val},i}^{t}-L_{\text{val},i}^{t+1}(p^{t,2})\\ \vdots\\ L_{\text{val},i}^{t}-L_{\text{val},i}^{t+1}(p^{t,m})\end{bmatrix}$ (14)
Let $P\in\mathbb{R}^{m\times m}$ be the leftmost matrix and $\beta_{i}\in\mathbb{R}^{m}$ be the vector on the right hand side. Then, we can write $A_{i}^{t}=P^{-1}\beta_{i}$. We solve this system for each $i\in[m]$ to obtain $A^{t}$.
The advantage of this method is that it directly estimates the optimal $A^{t\star}$ that is used in the mixing law. However, it requires $m$ sweeps per round, because the key quantity we must observe to learn $A^{t}$ is $L_{\text{val},i}^{t}-L_{\text{val},i}^{t+1}(\bm{p})$: the change in loss after training through the entire round $t$. As a result, this approach requires $m$ extra full training runs to learn $A^{t}$. Below, we will describe how we can compute cheaper alternatives to $L_{\text{val},i}^{t}-L_{\text{val},i}^{t+1}(\bm{p})$.
D.2 Modification 1: shortening training sweeps
This modification is depicted in Figure 7 (middle). A simple way to reduce the number of extra training runs needed to estimate $A^{t}$ is to train on each mixture $p^{t,j}$ for less than a round. Let $\delta$ denote the fraction of the round we use for the training sweep. Then, our system of equations in 14 uses $L_{\text{val},i}^{t}-L_{\text{val},i}^{t+\delta}(p^{t,j})$; we simply record the loss difference over $\delta$ of the round rather than the entire round, and use this to solve for $A^{t}$. Now, this approach effectively requires $m\delta$ extra training runs; however, this cost is still linear in the number of data groups. Moreover, there is some inaccuracy incurred by using $\delta$ of a round to approximate the entire round.
D.3 Modification 2: “interleaving” training sweeps
This modification is depicted in Figure 7 (bottom). Our final modification to derive LearnParams is to convert the training sweep—where we checkpoint the model and execute $m$ separate runs for $\delta$ of a round—into one round without requiring any checkpointing or rolling back of training. Our intuition is that if we interleave different mixtures sequentially at a high frequency, we can simulate executing these mixtures simultaneously. This is similar to a concept in signal processing called time-division multiplexing, in which two or more signals or bit streams are transferred appearing simultaneously as sub-channels in one communication channel, but are physically taking turns on the channel [^1] https://en.m.wikipedia.org/wiki/Time-division_multiplexing.
Formally, we break down the $\delta S/T$ steps allocated for learning $A^{t}$ into $K$ intervals, where $K=mk$ and $k$ is the number of sweeps per mixture. We construct an interleaved order of $p^{t,1},\dots,p^{t,m}$ over these $K$ intervals, and we denote their index order as $\mathcal{I}\in[m]^{K}$. Let $\mathcal{I}_{\tau}$ denote the mixture at the $\tau$th position in $\mathcal{I}$. We can denote the model at the end of each interval as $t+\delta/K,t+2\delta/K,\dots,t+\delta$. During the $\tau$th interval, we train on one $p^{t,\mathcal{I}_{\tau}}$ and observe the change in loss, $L_{\text{val},i}(f^{t+(\tau-1)\delta/K})-L_{\text{val},i}(f^{t+\tau\delta/K})$ for each validation group $i$. Let $\mathcal{T}_{j}=\{\tau:\mathcal{I}_{\tau}=j\}$ be all the intervals where $p^{t,j}$ is assigned. We approximate $L_{\text{val},i}^{t}-L_{\text{val},i}^{t+1}(p^{t,j})$ with $\frac{1}{|\mathcal{T}_{j}|}\sum_{\tau\in\mathcal{T}_{j}}^{k}L_{\text{val},i}(f^{t+(\tau-1)\delta/K})-L_{\text{val},i}(f^{t+\tau\delta/K})$. These approximated loss differences are then used to recover $A^{t}$ from the system of linear equations.
Lastly, note that the choice of $k$ controls the interleaving frequency and the bias of the estimated $A^{t}$. Suppose that $k=1$. This means that each mixture is only assigned to one interval, and this could be at the beginning, middle, or end of the $\delta$ round. Then, the change in loss is a poor approximation of the original quantity $L_{\text{val},i}^{t}-L_{\text{val},i}^{t+1}(\bm{p})$ due to dependence on time. However, as we increase $k$, the mixture $p^{t,j}$ will be trained on in the beginning, middle, and end of the $\delta$ round, allowing for a less time-biased estimate of the loss change.
With this modification, LearnParams now requires no extra training. However, there are still some performance tradeoffs. First, in order to save compute, our estimate of $A^{t}$ via the shortened interleaved sweeps is less accurate than the naive approach. Second, without rolling back training, Aioli has both an “explore” and “exploit” phase, where the former learns $A^{t}$ over $\delta$ of the round and the latter uses $A^{t}$ to set $p^{t}$ and mix data accordingly for the remainder of the round. If $\delta$ is large, the estimate of $A^{t}$ may be relatively more accurate. However, training for longer on the sweep mixtures $p^{t,1},\dots p^{t,m}$ may be suboptimal for the performance of the model. Moreover, the training duration that utilizes the $p^{t}$ that is updated using the more accurate $A^{t}$ is now shortened. Therefore, adjusting $\delta$ is key to ensuring that $A^{t}$ is accurate and the model performs well.
Appendix E Experimental Details
E.1 Data
To obtain a test set, we shuffle and split the validation set from SlimPajama-6B [54, 73] in half.
[…]
E.1.1 Training
@@
Here, we discuss the training setups for the restricted and unrestricted settings. For the $m=2,3$ settings, we train a 160M model using Pythia-160M’s configuration for $S=5000$ steps and results are averaged over 5 random seeds. For $m=7$, we train a 160M model using Pythia-160M’s configuration for $S=40000$ steps results are averaged over 3 random seeds. All settings use FlashAttention [18], batch size of 8, context size of 2048, and cosine learning rate decay from a starting learning rate of 5e-5 to 1e-5 with 500 steps of learning rate warmup.
@@
E.2 Data mixing methods
@@
In the unrestricted setting, we found it sometimes helpful to use an exponential moving average with proportion $\gamma$ over $A^{t}$ for Aioli. Formally, the standard $p^{t}$ update rule in Algorithm 1 can be unrolled as $p_{j}^{t+1}\propto p_{j}^{0}\exp(\eta\sum_{\tau=1}^{t}\sum_{i=1}^{m}A_{ij}^{\tau})$, which places equal weight on every $A_{ij}^{\tau}$. To incorporate the EMA, we define $A^{1}_{\text{ema}}=\bar{A}^{1}$ and $A^{t}_{\text{ema}}=(1-\gamma)\bar{A}^{t}+\gamma A^{t-1}_{\text{ema}}$. We then use the update rule $p_{j}^{t+1}\propto p_{j}^{0}\exp(\eta A_{\text{ema}}^{t})$. This allows Aioli to gradually decay the contributions of $A^{t}$, such that the value of $p^{t}$ is less dependent on earlier proportions in the training.
We summarize the hyperparameters used in Aioli, providing their default values as well as guidelines for how to set them. Refer to Algorithm 1 and 2 to see how they are used:
  • Number of rounds $T$: we set this to $20$ in all experiments. Larger $T$ means more frequent updates to the mixture proportions.
  • Sweeps $k$: we set this to be $4$ for $m=2,3$ and $2$ for the full SlimPajama experiments. We did not adjust this hyperparameter otherwise. Intuitively, a larger $k$ will give a more accurate $A^{t}$, because this means that each $p^{i,t}$ will be trained on more frequently throughout the $\delta$ proportion of the round; however, this will also result in less of the round being allocated to exploiting $A^{t}$ via using $p^{t}$.
  • $\varepsilon$ one-hot smoothing factor: we set this to be $0.75$ in all experiments. In general, $\varepsilon$ must be set between 00 and $1$, where 00 results in the training sweep using one-hot mixture proportions to learn $A^{t}$, which means that each batch only consists of one data group and can result in poor learning dynamics. $\varepsilon=1$, on the other hand, means that our training sweep would only consist of uniform proportions.
  • EGD step size $\eta$: we sweep $\{0.1,~{}0.2,~{}0.3,~{}0.5\}$, with higher $\eta$ resulting in greater magnitude of the proportion update.
  • Proportion of round $\delta$ dedicated to learning $A^{t}$: We use $\delta=0.128,0.288,0.007$ for $m=2,3,7$, respectively. Intuitively, a larger $\delta$ will give more accurate $A^{t}$ because the parameter is learned on more data, but this will also result in less of the round being allocated to exploiting $A^{t}$ via using $p^{t}$.
  • EMA parameter $\gamma$: we sweep None, $0.1,0.5$. Intuitively, None means that the $p^{t}$ update is equally dependent on all previous $p^{t}$’s, while a small $\gamma=0$ means that the $p^{t}$ update is only a function of the current $A^{t}$.
For the last three hyperparameters, $\eta,\delta,\gamma$, we used different values of them in different experiments. Tables 1213141516, and 17 list exact values for the unrestricted and restricted settings for $m=2,3,7$. In addition, Appendix F.3 provides results on hyperparameter sensitivity for $\eta,\delta$, and $\gamma$.
Data groups Hyperparameter Value
arXiv/SE $\cdot$ proportion of round $\delta$ 0.128
$\cdot$ EGD learning rate $\eta$ 0.2
$\cdot$ EMA parameter $\gamma$ 0.1
GitHub/C4 $\cdot$ proportion of round $\delta$ 0.128
$\cdot$ EGD learning rate $\eta$ 0.3
$\cdot$ EMA parameter $\gamma$ 0.5
Books/SE $\cdot$ proportion of round $\delta$ 0.128
$\cdot$ EGD learning rate $\eta$ 0.1
$\cdot$ EMA parameter $\gamma$ None
@@
Data groups Hyperparameter Value
arXiv/SE $\cdot$ proportion of round $\delta$ 0.128
@@
GitHub/C4 $\cdot$ proportion of round $\delta$ 0.128
@@
Books/SE $\cdot$ proportion of round $\delta$ 0.128
@@
Data groups Hyperparameter Value
arXiv/Books/SE $\cdot$ proportion of round $\delta$ 0.288
$\cdot$ EGD learning rate $\eta$ 0.5
$\cdot$ EMA parameter $\gamma$ None
CommonCrawl/GitHub/Wiki $\cdot$ proportion of round $\delta$ 0.288
$\cdot$ EGD learning rate $\eta$ 0.3
$\cdot$ EMA parameter $\gamma$ 0.5
Data groups Hyperparameter Value
arXiv/Books/SE $\cdot$ proportion of round $\delta$ 0.288
@@
$\cdot$ EMA parameter $\gamma$ None
CommonCrawl/GitHub/Wiki $\cdot$ proportion of round $\delta$ 0.288
$\cdot$ EGD learning rate $\eta$ 0.2
$\cdot$ EMA parameter $\gamma$ None
Data groups Hyperparameter Value
SlimPajama, full $\cdot$ proportion of round $\delta$ 0.07
@@
Data groups Hyperparameter Value
SlimPajama, full $\cdot$ proportion of round $\delta$ 0.07